# Three dimensional end-to-end simulation for kilonova emission from a black-hole neutron-star merger

### v1

@@ 3 The BH-NS model
In this work, we employ the NR outflow profiles and nucleosynthetic data obtained in Hayashi et al. (2022, 2023) and Wanajo et al. (2022) as the input for the HD simulations. In particular, we employ the outflow data of model Q4B5H in Hayashi et al. (2022). For this model, a BH-NS binary of which the NS mass, BH mass and dimensionless spin are initially $1.35\,M_{\odot}$, $5.4\,M_{\odot}$ (thus 4 times larger than the NS mass), and 0.75, respectively, is considered with the DD2 EOS (Banik et al., 2014). The poloidal magnetic field with the maximum strength of $5\times 10^{16}\,{\rm G}$ is initially set in the NS, while the resulting ejecta profile is not sensitive to the initial magnetic-field strength and configuration (Hayashi et al., 2023). We set $6.6\,M_{\odot}$ as the BH mass of the metric employed in the HD simulations, which approximately agrees with the summation of the remnant BH mass and matter outside the BH measured at $t=0.1\,{\rm s}$.
For model Q4B5H, the NS experiences significant tidal disruption before it reaches the inner-most stable circular orbit of the binary ($t\approx 10\,{\rm ms}$). This leads to the formation of massive ejecta and torus around the remnant BH. Ejecta formed at the time of the NS tidal disruption, which often referred to as the dynamical ejecta, are concentrated in the vicinity of equatorial plane and exhibit significant non-axisymmetric geometry. The dynamical ejecta typically have low electron fraction (0.03–0.07) because those are driven primarily by gravitational torque and do not go through significant weak processes in the merger.
@@
In this paper, we study the ejecta and KN property for one case of a BH-NS merger among available NR results as the first step for the end-to-end kilonova simulation. However, we should note that the disk and ejecta masses formed in BH-NS mergers can have large variety depending on the binary parameters, such as the BH and NS masses, BH spin, and NS radius (Rosswog, 2005; Shibata & Taniguchi, 2008; Etienne et al., 2009; Lovelace et al., 2013; Kyutoku et al., 2015; Foucart et al., 2018), as well as the adopted EOS (Hayashi et al., 2023). For example, the smaller amount of disk and ejecta would be formed for the case that the NS radius is smaller due to softer EOS, such as the SFHo EOS (Steiner et al., 2013; Hayashi et al., 2023). Hence, the resulting property of the KN light curves can also have a large diversity. Therefore, we emphasize that the ejecta and KN property found for model Q4B5H with the DD2 EOS may not be universal property for every case of BH-NS mergers, and we leave the investigation of the binary parameter and EOS dependences for a future work.
@@ 4 Results: Hydrodynamics simulation
4.1 Ejecta mass evolution
Figure 1: Time evolution of the total rest mass in the computational domain of the HD simulation (the purple curve). The green curve denotes the same as for the purple curve but only for the matter which satisfies the geodesic criterion ($u_{t}<-1$ where $u_{t}$ is the lower time component of the four velocity). The blue dashed curve with the label “input" denotes the rest mass obtained by integrating the mass flux of the NR outflow data which is employed as the inner boundary condition of the HD simulation. The black dashed line denotes the time at which the NR outflow data run out.
@@ 4.1 Ejecta mass evolution
After the NR outflow data run out at $t\approx 1\,{\rm s}$, we impose a floor rest-mass density value to the inner boundary. It is clearly seen in Figure 1 that the total mass in the computational domain decreases after that time, indicating that the matter is artificially falling back and escaping through the inner boundary. This happens because the pressure support from the inner boundary vanishes after the outflow data run out. However, the total mass of the matter with gravitationally unbound orbits remains increasing even after the time when the NR outflow data run out, as the consequence of the acceleration of the matter in the presence of the thermal pressure gradient.
@@ 4.2 Ejecta profiles at homologously expanding phase
Figure 2: Rest-mass density and electron fraction ($Y_{e}$) profiles on the equatorial plane at $t=0.1\,{\rm d}$. The yellow dotted lines denote the angles for which the meridional ejecta profiles are shown in Figures 3 and 4. The white dotted curves denote the longitudinal angle ranges in which the KN light curves shown in Figures 8 and  9 are obtained. The value employed as the initial condition of the nucleosynthesis calculation is shown in the $Y_{e}$ profile (see Appendix B and Wanajo et al. (2022) for the detail.)
@@
On the other hand, the post-merger ejecta exhibit approximately an axisymmetric shape. It has two distinct components with one having approximately a spherical shape and the other having the cone-like shape. The former is concentrated in the region within $\approx 0.05ct$ while the latter is more extended in the vertical direction with the polar opening angle of $\approx 10^{\circ}$ and the vertical extent reaches $\approx 0.25ct$. As we show below, this complex geometry of the post-merger ejecta is realized by the interaction with the dynamical ejecta which significantly expand due to thermal pressure enhanced by radioactive heating.
@@
Figure 5 shows the rest-mass density and electron fraction profiles of ejecta on the equatorial and meridional planes at $t=0.1\,{\rm d}$ obtained by the HD simulation but switching off radioactive heating. Under the presence of radioactive heating, the dynamical ejecta expand significantly due to the increase in thermal pressure and the inhomogeneities in the rest-mass density are also smoothed out, as clearly seen in Figures 2 and 3. These results are consistent with the finding of Rosswog et al. (2014); Grossman et al. (2014) in the context of BNSs, and of Fernández et al. (2015); Darbha et al. (2021) in the context of BH-NSs. In fact, the resulting aspect ratio of the dynamical ejecta is found to be close to unity as the model H4 in Darbha et al. (2021). The radioactive heating rate of the dynamical ejecta in our model also agrees with that in Darbha et al. (2021) (see Figure 6).
@@
The comparison between Figure 4 and Figure 5 shows that the profile of the post-merger ejecta is affected by the modification of the dynamical eject profile. Figure 5 shows that, in the absence of radioactive heating, the post-merger ejecta exhibit a prolate shape with the extension of $0.1ct$ and $0.25ct$ in the equatorial and vertical directions, respectively. On the other hand, the radioactive heating significantly expands the dynamical ejecta, which compress the post-merger ejecta in $0.05ct\leq z\leq 0.15ct$ and confine the ejecta in the region of $\lesssim 0.05ct$ as found in Figure 4. This happens because of the higher typical electron fraction of the post-merger ejecta. The higher electron fraction leads to the relatively small radioactive heating rate and hence small enhancement of the pressure of the ejecta compared to the dynamical component.
@@
Significant expansion of the dynamical ejecta and enforced confinement of the post-merger ejecta in the presence of radioactive heating are not found in the BNS models in our previous studies (Kawaguchi et al., 2021, 2022, 2023). This is because the dynamical ejecta of the present BH-NS model and the BNS models studied in Rosswog et al. (2014); Grossman et al. (2014) are massive compared to the post-merger ejecta and also much more confined around the equatorial plane compared to the BNS models in our previous studies. As a result, higher internal energy density and high thermal pressure are realized. Hence, the importance of the radioactive heating will depend on the density and isotopic-abundance profiles of ejecta, which can have a variety even among BH-NS mergers depending on the binary parameters or the adopted EOS.
@@ 5 Results: KN light curves
Figure 7: Total bolometric luminosity and total energy deposition rate for model Q4B5H. The isotropically equivalent bolometric luminosity observed in AT2017gfo with the distance of 40 Mpc is shown by the filled circles adopting the data in Waxman et al. (2018), which assume a black-body fit to the photo-metric observations. Note that the bolometric light curve before 1 d is hidden since it is not reliable due to the lack of opacity data in the high temperature regime ($\gtrsim 20,000\,{\rm K}$).
@@
Figure 7 shows the bolometric luminosity calculated by the RT simulation employing the ejecta rest-mass density, elemental abundance, and radioactive heating rate profiles obtained by the combination of the results of the HD simulation and nucleosynthesis calculation (Wanajo et al., 2022). The total energy deposition rate taking the thermalization efficiency into account is also plotted in Figure 7. As we mentioned in Section 2, our atomic data include only up to the triple ionization for all the ions, and the opacity of ejecta in the early phase ($t\leq 0.5\,{\rm d}$) may be underestimated due to high temperature ($\gtrsim 20,000\,{\rm K}$). Hence, hereafter, we only focus on the light curves after 1 d.
@@
For 1–10 d, the bolometric luminosity is approximately constant with the value of 1–2×$\times 10^{\rm 41}\,{\rm erg/s}$. It decreases only slowly and the change is only by a factor of 2 during this epoch. However, after 10 d, the bolometric luminosity starts decreasing more rapidly, and it decreases by a factor of 5 during 10–30 d. This faint and long-lasting emission is caused by the fact that it is primarily powered by the lanthanide-rich dynamical ejecta, in which a long photon diffusion time scale is realized by the large mass and high opacity. This behaviour of the bolometric luminosity is qualitatively the same as that found in the models with massive dynamical ejecta studied in the previous study (MS1Q3a75 and H4Q3a75 in Tanaka et al. (2014)). The bolometric luminosity converges to the total deposition rate after 20 d, which suggests that the entire thermal photons created in the ejecta immediately diffuse out from the ejecta after this epoch. As we show below, however, the viewing-angle dependence of the emission due to the aspherical profile of the ejecta opacity is still playing a role up to 30 d.
@@
Our present BH-NS KN model shows significantly distinct light curves from the observation of AT2017gfo. Specifically, our BH-NS model is fainter by a factor of 2 around $\sim 1\,{\rm d}$ than AT2017gfo. However, due to the slow decrease in the bolometric luminosity, compared to AT2017gfo, our BH-NS KN model becomes comparably bright at 4 d, and brighter by a factor of 5 at 10 d. This result clearly shows that a BH-NS binary which we study particularly in this work is not likely to be the progenitor of AT2017gfo.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Figure 8: Isotropically equivalent bolometric luminosities observed from various viewing angles for model Q4B5H. The top panels denote the comparison among the results for different longitudinal directions, while the middle and bottom panels denote the comparison among the results for different latitudinal directions. The isotropically equivalent bolometric luminosity observed in AT2017gfo with the distance of 40 Mpc is also shown by the filled circles adopting the data in Waxman et al. (2018), which assume a black-body fit to the photo-metric observations. Note that the light curves before 1 d are hidden since they are not reliable due to the lack of opacity data in the high temperature regime ($\gtrsim 20,000\,{\rm K}$).
@@
Figure 8 shows the results of the isotropically equivalent bolometric light curves observed from various viewing angles for the present model. Focusing on the observer in the polar direction with $\theta\leq 28^{\circ}$, the KN emission is brightest in b): $135^{\circ}\leq\phi<180^{\circ}$. This direction approximately matches to the longitudinal direction in which the dynamical ejecta have the most of their mass (see Figures 2 and 3). On the other hand, the faintest emission is observed from the direction in which the dynamical ejecta least present (‘d)’: $315^{\circ}\leq\phi<360^{\circ}$). Nevertheless, the variation in the bolometric luminosity is not so large and it is always within 40%. This is reasonable because the observers with different longitudinal angles are in similar directions for the polar view $\theta\leq 28^{\circ}$.
@@
On the other hand, the longitudinal variation is larger for the equatorial view ($82^{\circ}\leq\theta\leq 90^{\circ}$). For this case, the KN emission is also brightest in b): $135^{\circ}\leq\phi<180^{\circ}$ in which direction the dynamical ejecta is most present. By contrast, the bolometric luminosity is the faintest in a): $45^{\circ}\leq\phi<90^{\circ}$. This is because, in this direction, the relatively thin part of the dynamical ejecta present in $R\gtrsim 0.3ct$ around the equatorial plane (see Figure 3) suppresses radiation from the ejecta center (note that such suppression is not significant from the direction of c): $225^{\circ}\leq\phi<270^{\circ}$ due to the absence of the low density ejecta above $R\approx 0.3ct$). Meanwhile, the emission from the post-merger ejecta enhances the luminosity in this view ($82^{\circ}\leq\theta\leq 90^{\circ}$) in which the dynamical ejecta are least present (‘d)’: $315^{\circ}\leq\phi<360^{\circ}$). The variation in the bolometric luminosity is larger for the equatorial view than that for the polar view, and it is larger than a factor of 2 for 1–7 d.
@@
Interestingly, the latitudinal viewing-angle dependence is still present in d): $315^{\circ}\leq\phi<360^{\circ}$ even after 20 d at which the total bolometric luminosity converges to the total deposition rate (see Figure 7). This can be understood by the fact that the post-merger ejecta are present in between the high-density part of the dynamical ejecta and the observer in the direction of d): $315^{\circ}\leq\phi<360^{\circ}$ (see Figures 2 and  3). While the post-merger ejecta have a minor contribution to the luminosity in the late epoch due to their relatively low heating rate, they still contribute as an opacity source (because of the relatively high density) to prevent photons emitted in the high-density part of the dynamical ejecta diffusing out to the direction of d): $315^{\circ}\leq\phi<360^{\circ}$. Although it is not quantitatively significant, this long-lasting viewing-angle dependence due to the non-axisymmetric geometry of the ejecta might have an impact on estimating the total deposition rate in the ejecta from the late-time observation.
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 9: gzK-band light curves for model Q4B5H observed from various viewing angles with the distance of 40 Mpc. The top panels denote the comparison among the results for different longitudinal directions, while the middle and bottom panels denote the comparison among the results for different latitudinal directions. The data points denote the AB magnitudes of AT2017gfo taken from Villar et al. (2017).
Figure 9 shows the optical (the g and z bands) and near-infrared (the K-band) light curves observed from various viewing angles. As is the case for the bolometric luminosity, for the polar view ($\theta\leq 28^{\circ}$), the gzK-bands are the brightest and the faintest in b): $135^{\circ}\leq\phi<180^{\circ}$ and in d): $315^{\circ}\leq\phi\leq 360^{\circ}$, in which the dynamical ejecta are mostly and least present, respectively. The viewing-angle dependence of the emission is weak from the polar view, and the variation is always within 0.5 mag around the peak magnitudes.
For the equatorial view ($\theta\geq 28^{\circ}$), the gzK-band emission is the brightest in b): $135^{\circ}\leq\phi<180^{\circ}$, while the emission in a): $45^{\circ}\leq\phi\leq 90^{\circ}$ becomes the faintest. The longitudinal viewing-angle dependence of the emission is significant in the gz-bands, and it is always larger than 1 mag among different longitudinal directions. The variation in the K-band magnitude among different longitudinal directions is relatively small compared to that in the gz-bands, and it is always approximately within 1 mag among the all viewing angles.

### v2

@@ 3 The BH-NS model
In this work, we employ the NR outflow profiles and nucleosynthetic data obtained in Hayashi et al. (2022, 2023) and Wanajo et al. (2022) as the input for the HD simulations. In particular, we employ the outflow data of model Q4B5H in Hayashi et al. (2022). For this model, a BH-NS binary of which the NS mass, BH mass and dimensionless spin are initially $1.35\,M_{\odot}$, $5.4\,M_{\odot}$ (thus 4 times larger than the NS mass), and 0.75, respectively, is considered with the DD2 EOS (Banik et al., 2014). The poloidal magnetic field with the maximum strength of $5\times 10^{16}\,{\rm G}$ is initially set in the NS; with this setting, the magnetorotational instability (Balbus & Hawley, 1998) in the remnant disk is resolved and the $\alpha$-$\Omega$ dynamo subsequently in action can be captured. We set $6.6\,M_{\odot}$ as the BH mass of the metric employed in the HD simulations, which approximately agrees with the summation of the remnant BH mass and matter outside the BH measured at $t=0.1\,{\rm s}$.
We note that, although BHs may be born with low spins (e.g., Fuller & Ma, 2019; Gottlieb et al., 2023a), we assume a relatively large spin (dimensionless BH spin of 0.75), which tends to give a larger ejecta mass and brighter kilonova  (e.g., Colombo et al., 2024). We also note that the large initial magnetic field is artificially set for the numerical prescription, while Hayashi et al. (2023) showed that the resulting ejecta profile may not be very sensitive to the initial magnetic-field strength and configuration. However it is important to keep in mind systematic errors associated with these uncertainty.
For model Q4B5H, the NS experiences tidal disruption before it reaches the inner-most stable circular orbit of the binary ($t\approx 10\,{\rm ms}$). This leads to the formation of massive ejecta and disk around the remnant BH. Ejecta formed at the time of the NS tidal disruption, which often referred to as the dynamical ejecta, are concentrated in the vicinity of equatorial plane and exhibit significant non-axisymmetric geometry. The dynamical ejecta typically have low electron fraction (0.03–0.07) because those are driven primarily by gravitational torque and do not appreciably go through weak processes in the merger.
@@
In this paper, we study the ejecta and KN property for one case of a BH-NS merger among available NR results as the first step for the end-to-end KN simulation. However, we should note that the disk and ejecta masses formed in BH-NS mergers can have large variety depending on the binary parameters, such as the BH and NS masses, BH spin, and NS radius (Rosswog, 2005; Shibata & Taniguchi, 2008; Etienne et al., 2009; Lovelace et al., 2013; Kyutoku et al., 2015; Foucart et al., 2018), as well as the adopted EOS (Hayashi et al., 2023). For example, the smaller amount of disk and ejecta would be formed for the case that the NS radius is smaller due to softer EOS, such as the SFHo EOS (Steiner et al., 2013; Hayashi et al., 2023). Hence, the resulting property of the KN light curves can also have a large diversity. Therefore, we emphasize that the ejecta and KN property found for a particular model (Q4B5H) with the DD2 EOS may not be universal property for every case of BH-NS mergers, and we leave the investigation of the binary parameter and EOS dependences for future work.
@@ 4 Results: Hydrodynamics simulation
4.1 Ejecta mass and energy evolution
Figure 1: Time evolution of the total rest mass in the computational domain of the HD simulation (the purple curves). The solid and dotted purple curves denote the results for the HD simulations in which radioactive heating are turned on and off, respectively. The green curve denotes the same as for the purple curve but only for the matter which satisfies the geodesic criterion ($u_{t}<-1$ where $u_{t}$ is the lower time component of the four velocity). The blue dashed curve with the label “input" denotes the rest mass obtained by integrating the mass flux of the NR outflow data which is employed as the inner boundary condition of the HD simulation. The black dashed line denotes the time at which the NR outflow data run out.
@@ 4.1 Ejecta mass and energy evolution
After the NR outflow data run out at $t\approx 1\,{\rm s}$, we impose a floor rest-mass density value to the inner boundary. It is clearly seen in Figure 1 that the total mass in the computational domain decreases after that time, indicating that the matter is artificially falling back and escaping through the inner boundary. This happens because the pressure support from the inner boundary vanishes after the outflow data run out. However, the total mass of the matter with gravitationally unbound orbits remains increasing even after the time when the NR outflow data run out, as a consequence of the acceleration of the matter in the presence of the thermal pressure gradient. In fact, the mass of the fall back matter is larger due to lower pressure for the case in which radioactive heating is turned off (see the purple dotted curve in Figure 1).
@@
Figure 2: Time evolution of the ejecta kinetic energy and total released nuclear binding energy. The top and bottom panels show the total value and mass-averaged value per baryon, respectively. The black solid and dotted curves denote the kinetic energy for the HD simulations in which radioactive heating are turned on and off, respectively. The purple curve denotes the total nuclear binding energy released during the radioactive decay. The green and blue curves denote the same as the purple curve but only taking into account the ejecta components which pass the inner boundary at $t<0.3\,{\rm s}$ (the dynamical ejecta) and $t\geq 0.3\,{\rm s}$ (the post-merger ejecta), respectively.
Figure 2 shows the time evolution of the ejecta kinetic energy and total released nuclear binding energy. The top panel of Figure 2 shows that the ejecta kinetic energy increases around $t\sim 0.04\,{\rm s}$ and $t\sim 0.5\,{\rm s}$. The former increase is due to the entry of the dynamical ejecta into the computational domain. On the other hand, the increase at $t\sim 0.5\,{\rm s}$ is due to the acceleration induced by the radioactive heating. In fact, the increase is not found for the case in which the radioactive heating is turned off. The increase in the kinetic energy by the radioactive heating agrees with $\approx 70\%$ of the released nuclear binding energy; the 30% reduction is owing to the neutrino loss. The slight decrease in the kinetic energy after $\sim 0.1\,{\rm s}$ is due to the fall back of the matter to the outside of the computational domain. After $t\sim 1000\,{\rm s}$ the kinetic energy becomes mostly a constant, which indicates that the homologous expansion of the system is achieved. The radioactive heating is dominated by the component which passes the inner boundary at $t<0.3\,{\rm s}$, that is, the dynamical ejecta.
The bottom panel of Figure 2 shows the mass-averaged values of the ejecta kinetic energy and total released nuclear binding energy per baryon. The released nuclear binding energy saturates with $\approx 5\,{\rm MeV}/{\rm baryon}$. That for the dynamical ejecta saturates with a larger value ($\approx 6\,{\rm MeV}/{\rm baryon}$) than the post-merger ejecta ($\approx 2\,{\rm MeV}/{\rm baryon}$). This indicates that the radioactive heating in the dynamical ejecta is more efficient because of the stronger $r$-process with fission recycling. Note that the difference in the mass-averaged kinetic energy per baryon between the HD simulations in which radioactive heating are turned on and off is less pronounced compared to that in the top panel of Figure 2. This is because the mass of the fall-back matter is larger for the case without radioactive heating (see Figure 1), and hence, the mass-averaged value increases.
@@ 4.2 Ejecta profiles at homologously expanding phase
Figure 3: Rest-mass density and electron fraction ($Y_{e}$) profiles on the equatorial plane at $t=0.1\,{\rm d}$. The yellow dotted lines denote the angles for which the meridional ejecta profiles are shown in Figures 4 and 5. The white dotted curves denote the longitudinal angle ranges in which the KN light curves shown in Figures 9 and 10 are obtained.
@@
On the other hand, the post-merger ejecta exhibit a more axisymmetric shape. It has two distinct components with one having approximately a spherical shape and the other having the cone-like shape. The former is concentrated in the region within $\approx 0.05ct$ while the latter is more extended in the vertical direction with the polar opening angle of $\approx 10^{\circ}$ and the vertical extent reaches $\approx 0.25ct$. As we show below, this complex geometry of the post-merger ejecta is realized by the interaction with the dynamical ejecta which significantly expand due to thermal pressure enhanced by radioactive heating.
@@
Figure 6 shows the rest-mass density and electron fraction profiles of ejecta on the equatorial and meridional planes at $t=0.1\,{\rm d}$ obtained by the HD simulation but switching off radioactive heating. Under the presence of radioactive heating, the dynamical ejecta expand significantly due to the increase in thermal pressure and the inhomogeneities in the rest-mass density are also smoothed out, as clearly seen in Figures 3 and 4. These results are consistent with the finding of Rosswog et al. (2014); Grossman et al. (2014) in the context of BNSs, and of Fernández et al. (2015); Darbha et al. (2021) in the context of BH-NSs. In fact, the resulting aspect ratio of the dynamical ejecta is found to be close to unity as the model H4 in Darbha et al. (2021). The radioactive heating rate as a function of time for the dynamical ejecta in our model is similar to that in Darbha et al. (2021) (see Figure 7) , although the latter appears to be a factor of a few larger than ours [^2] We note that, although it is claimed to be realistic heating rates for $Y_{e}\approx 0.1$ (according to Metzger et al. 2010; Lippuner & Roberts 2015), the heating rate model H4 in Darbha et al. (2021) is unphysical such that the total amount of energy released in this model is $\approx 13$ MeV/baryon (larger than the highest binding energy among radioactively stable nuclei, 8.8 MeV/baryon for 62Ni). On the other hand, the total amount of energy released by radioactive heating in our model is $\approx 5$ MeV/baryon (see Figure 2), which is consistent with the typical value $\approx 8$ MeV/baryon for stable $r$-process nuclei by considering the neutrino-energy loss..
@@
The comparison between Figures 5 and 6 shows that the profile of the post-merger ejecta is modified by the dynamical ejecta. Figure 6 shows that, in the absence of radioactive heating, the post-merger ejecta exhibit a prolate shape with the extension of $0.1ct$ and $0.25ct$ in the equatorial and vertical directions, respectively. On the other hand, the radioactive heating significantly expands the dynamical ejecta, which compress the post-merger ejecta in $0.05ct\leq z\leq 0.15ct$ and confine the ejecta in the region of $\lesssim 0.05ct$ as found in Figure 5. This is due to the higher typical electron fraction of the post-merger ejecta. The higher electron fraction results in the relatively small radioactive heating rate and hence small enhancement of the pressure of the ejecta compared to the dynamical component (see Figure 2).
We note that, while the dynamical ejecta apparently expand significantly, the total increase in the ejecta total kinetic energy is not as significant as that one might expect from the change in the ejecta edge ($\approx 0.1\,c\rightarrow\approx 0.2\,c$ in the $z$-direction and $\approx 0.3\,c\rightarrow\approx 0.4\,c$ in the $R$-direction). This is because only a small fraction of the ejecta matter is accelerated to such high velocity. We stress that the increase in the ejecta kinetic energy is consistent with the energy deposited by radioactive heating (see Figure 2).
@@
Significant expansion of the dynamical ejecta and enforced confinement of the post-merger ejecta in the presence of radioactive heating are not found in the BNS models in our previous studies (Kawaguchi et al., 2021, 2022, 2023). This is because the dynamical ejecta of the present BH-NS model, as well as of the BNS models studied in Rosswog et al. (2014); Grossman et al. (2014), are more massive than the post-merger ejecta, and in addition, much more confined around the equatorial plane than in the BNS models. As a result, higher internal energy density and high thermal pressure are realized. This illustrates that the importance of the radioactive heating will depend on the density and isotopic-abundance profiles of ejecta, which can have a variety even among BH-NS mergers depending on the binary parameters or the adopted EOS.
@@ 5 Results: KN light curves
Figure 8: Total bolometric luminosity (solid curve) and total energy deposition rate (dashed curve). The isotropically equivalent bolometric luminosity observed in AT2017gfo with the distance of 40 Mpc is shown by the filled circles adopting the data in Waxman et al. (2018), which assume a black-body fit to the photo-metric observations. Note that the bolometric light curve before 1 d is hidden since it is not reliable due to the lack of opacity data in the high temperature regime ($\gtrsim 20,000\,{\rm K}$).
@@
Figure 8 shows the bolometric luminosity calculated by the RT simulation employing the profiles of the ejecta rest-mass density, elemental abundance, and radioactive heating rate obtained by the combination of the results of the HD simulation and nucleosynthesis calculation (Wanajo et al., 2022). The total energy deposition rate taking the thermalization efficiency into account is also plotted in Figure 8. As we mentioned in Section 2, our atomic data include only up to the triple ionization for all the ions, and the opacity of ejecta in the early phase ($t\leq 0.5\,{\rm d}$) may be underestimated due to high temperature ($\gtrsim 20,000\,{\rm K}$). Hence, hereafter, we only focus on the light curves after 1 d.
@@
For 1–10 d, the bolometric luminosity is approximately constant with the value of 1–2×$\times 10^{\rm 41}\,{\rm erg/s}$. It decreases only slowly and the change is only by a factor of 2 during this epoch. However, after 10 d, the bolometric luminosity starts decreasing more rapidly, and it decreases by more than a factor of 3 during 10–30 d. This faint and long-lasting emission is caused by the fact that it is primarily powered by the lanthanide-rich dynamical ejecta, in which a long photon diffusion time scale is realized by the large mass and high opacity. This behaviour of the bolometric luminosity is qualitatively the same as that found in the models with massive dynamical ejecta studied in a pioneering study (MS1Q3a75 and H4Q3a75 in Tanaka et al. 2014). The bolometric luminosity converges to the total deposition rate after 20 d, which suggests that the entire thermal photons created in the ejecta immediately diffuse out from the ejecta after this epoch. As we show below, however, the viewing-angle dependence of the emission due to the aspherical profile of the ejecta opacity still plays a role up to 30 d.
@@
Our present BH-NS KN model shows significantly distinct light curves from the observation of AT2017gfo. Specifically, our BH-NS model is fainter by a factor of 3 around $\sim 1\,{\rm d}$ than AT2017gfo. However, due to the slow decrease in the bolometric luminosity, compared to AT2017gfo, our BH-NS KN model becomes comparably bright at 5 d, and brighter by a factor of 5 at 10 d. This result clearly shows that a BH-NS binary which we study particularly in this work is not likely to be the progenitor of AT2017gfo.
@@
Figure 9: Isotropically equivalent bolometric luminosities observed from various viewing angles. The top panels compare the results among different longitudinal directions, while the middle and bottom panels compare the results among different latitudinal directions. The isotropically equivalent bolometric luminosity observed in AT2017gfo with the distance of 40 Mpc is also shown by the filled circles adopting the data in Waxman et al. (2018), which assume a black-body fit to the photo-metric observations. Note that the light curves before 1 d are hidden since they are not reliable due to the lack of opacity data in the high temperature regime ($\gtrsim 20,000\,{\rm K}$).
@@
Figure 9 shows the results of the isotropically equivalent bolometric light curves observed from various viewing angles for the present model. Focusing on the observer in the polar direction with $\theta\leq 28^{\circ}$, the KN emission is brightest in b): $135^{\circ}\leq\phi<180^{\circ}$. This direction approximately matches to the longitudinal direction in which the dynamical ejecta have most of their mass (see Figures 3 and 4). On the other hand, the faintest emission is observed from the direction in which the dynamical ejecta least present (‘d)’: $315^{\circ}\leq\phi<360^{\circ}$). Nevertheless, the variation in the bolometric luminosity is not so large and it is always within 40%. This is reasonable because the observers with different longitudinal angles are in similar directions for the polar view $\theta\leq 28^{\circ}$.
@@
On the other hand, the longitudinal variation is larger for the equatorial view ($82^{\circ}\leq\theta\leq 90^{\circ}$). For this case, the KN emission is also brightest in b): $135^{\circ}\leq\phi<180^{\circ}$ in which direction the dynamical ejecta are most abundant. By contrast, the bolometric luminosity is faintest in a): $45^{\circ}\leq\phi<90^{\circ}$ for $t\leq 5\,{\rm d}$. This is because, in this direction, the relatively thin part of the dynamical ejecta present in $R\gtrsim 0.3ct$ around the equatorial plane (see Figure 4) suppresses radiation from the ejecta center [^3] Note that such suppression is not significant from the direction of c): $225^{\circ}\leq\phi<270^{\circ}$ due to the absence of the low density ejecta above $R\approx 0.3ct$. Meanwhile, the emission from the post-merger ejecta enhances the luminosity in this view ($82^{\circ}\leq\theta\leq 90^{\circ}$) in which the dynamical ejecta are least present (‘d)’: $315^{\circ}\leq\phi<360^{\circ}$). The variation in the bolometric luminosity is larger for the equatorial view than that for the polar view, and it is larger than a factor of 2 for 1–7 d.
@@
Interestingly, the latitudinal viewing-angle dependence is still present in d): $315^{\circ}\leq\phi<360^{\circ}$ even after 20 d at which the total bolometric luminosity converges to the total deposition rate (see Figure 8). This can be understood by the fact that the post-merger ejecta are present in between the high-density part of the dynamical ejecta and the observer in the direction of d): $315^{\circ}\leq\phi<360^{\circ}$ (see Figures 3 and  4). While the post-merger ejecta have a minor contribution to the luminosity in the late epoch due to their relatively low heating rate, they still contribute as an opacity source (because of the relatively high density) to prevent photons emitted in the high-density part of the dynamical ejecta diffusing out to the direction of d): $315^{\circ}\leq\phi<360^{\circ}$. Although it is not quantitatively significant, this long-lasting viewing-angle dependence due to the non-axisymmetric geometry of the ejecta might have an impact on estimating the total deposition rate in the ejecta from the late-time observation.
@@
Figure 10: gzK-band light curves for observed from various viewing angles with the distance of 40 Mpc. The top panels compare the results among different longitudinal directions, while the middle and bottom panels compare the results among different latitudinal directions. The data points denote the AB magnitudes of AT2017gfo taken from Villar et al. (2017).
Figure 10 shows the optical (the g and z bands) and near-infrared (the K-band) light curves observed from various viewing angles. As is the case for the bolometric luminosity, for the polar view ($\theta\leq 28^{\circ}$), the gzK-bands are brightest and faintest in b): $135^{\circ}\leq\phi<180^{\circ}$ and in d): $315^{\circ}\leq\phi\leq 360^{\circ}$, in which the dynamical ejecta are mostly and least present, respectively. The viewing-angle dependence of the emission is weak from the polar view, and the variation is always within 0.5 mag around the peak magnitudes.
For the equatorial view ($82^{\circ}\leq\theta\leq 90^{\circ}$), the gzK-band emission is brightest in b): $135^{\circ}\leq\phi<180^{\circ}$, while the emission in a): $45^{\circ}\leq\phi\leq 90^{\circ}$ becomes faintest. The longitudinal viewing-angle dependence of the emission is significant in the gz-bands, and it is always larger than 1 mag among different longitudinal directions. The variation in the K-band magnitude among different longitudinal directions is relatively small compared to that in the gz-bands, and it is always approximately within 1 mag among the all viewing angles.